A disc (thin cylinder) of mass 0.1 kg and radius 0.41 m is rotating on a frictionless axle at a constant rate of 74 RPM.

What a the angular velocity and angular momentum of the disc?

A second disc of 0.05 kg and 0.18 m is dropped down the axle on top. The two then rotate together. What is the final angular velocity and RPM of the system of discs?
Answer #1

a) Angular velocity w = 74*2pi/60 = 7.749 rad/s answer

Angular momentum L = iw = 0.5*m*r^2*w = 0.5*0.1*0.41^2*7.749 = 0.06513 kg*m^2/s answer

b) By momentum conservation, i2*w2 = 0.06513

w2 = 0.06513/i2 = 0.06513/[0.5*0.1*0.41^2+0.5*0.05*0.18^2]

= 7.0678 rad/s answer

RPM = 7.0678*60/2pi = 67.49 ANSWER